OpenStudy (shayaan_mustafa):

In each part, a point is given in rectangular coordinates. Find two pairs of polar coordinates for the point, one pair satisfying r >= 0 and 0 <= θ < 2π , and the second pair satisfying r >= 0 and -π < θ <= π . (a) (-5,0) (b) (0,-2)

OpenStudy (shayaan_mustafa):

let clear question please.

OpenStudy (shayaan_mustafa):

me*

OpenStudy (experimentx):

point (1,2) in rectangular coordinates can be written in polar coordinates in two ways 1) calculate r \(r = \sqrt{x^2 + y^2} \) 2) calculate \( \theta \) \( \theta = \arctan{y/x}\) <--- theta can be negative or positive here now write it as \((r,\theta) \) or as \((-r,-\theta) \)

OpenStudy (shayaan_mustafa):

@experimentX I know this but don't know how to satisfy conditions friend.

OpenStudy (shayaan_mustafa):

a) r=5 and θ=0 rad Now what to do with conditions?

OpenStudy (experimentx):

what are a)s and b)'s ??

OpenStudy (shayaan_mustafa):

I have posted a)s and b)'s

OpenStudy (experimentx):

A is correct 0 <= 0

OpenStudy (shayaan_mustafa):

You mean r=5 and θ=0 rad is correct for part (a)?

OpenStudy (experimentx):

yea ...  Should have been like that ...

OpenStudy (shayaan_mustafa):

How can I satisfy conditions. I am just confusing about this dear. What to do for conditions?

OpenStudy (experimentx):

 You made it like that, which makes r negative.